Write the linear equation: perpendicular to x-3y=7 through (6,2). Thank you. :)
Answer by mananth(16946)

1 x + -3 y = 7
Find the slope of this line

-3 y = -1 x + 7
Divide by -3
y = 1/ 3 x + -2 1/3
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 1/3

The slope of a line perpendicular to the above line will be the negative reciprocal -3
Because m1*m2 =-1
The slope of the required line will be -3

m= -3 ,point ( 6 , 2 )
Find b by plugging the values of m & the point in
y=mx+b
2 = -18 + b
b= 20
m= -3
The required equation is y = -3 x + 20
m.ananth@hotmail.ca